# NeuroDeX: Unlocking Diverse Support in Decompiling Deep Neural Network Executables

### v1

@@ I Introduction
In recent years, deep learning (DL) has rapidly advanced in the real world. Deploying deep neural networks (DNNs) on edge devices can meet the real-time requirements of edge computing, enhance privacy protection and enable offline inference capabilities, making DNNs widely applicable in real-world scenarios  [2, 21, 45, 6]. DL compilers, such as TVM [3] and GLOW [37], can compile high-level DNN models into executables for inference on edge devices. DNNs are composed of different operators (e.g., Conv, Relu), and DL compilers compiles these operators into operator functions in executables. DL compilers optimize models during compilation to improve inference efficiency, which provides a good solution for deploying models on edge devices [55, 25, 22].
In DNN executables, the operators and weights are compiled into incomprehensible machine code, thereby reducing the risk of model extraction attacks compared to white-box deployment. However, DNN executables may still pose security risks due to decompilation. DNN executables can potentially be decompiled to original high-level models. This undermines the intellectual property of the model owners, especially for models trained on private data. Based on the recovered high-level models, attackers can perform white-box adversarial attacks and backdoor attacks, threatening the secure use of DNN executables.
Previous works [59, 53, 27, 42, 54] have attempted to decompile DNN binaries, but the existing methods have certain limitations. They struggle to simultaneously address compilation optimizations, support different architectures, and accommodate quantized compiled model, all of which are essential for meeting the demands of modern DL applications. These limitations can be summarized as follows. L1: The recognition for operator types is not accurate enough, previous methods struggle to handle fused operators under compilation optimization and rely on prior knowledge such as compiler versions. L2: The current cross-architecture decompilers rely on relatively heavy analysis methods or support too few types of operators. They also cannot analyze fused operators. L3: Previous works lack consideration for quantized models, which is a common inference optimization strategy for edge devices.
@@ II-A Deep Learning Compiler
DNNs are computational models consisting of multiple layers. These layers are constructed from different types of operators that perform various mathematical computations (e.g., Conv, Relu), and some operators have attributes (e.g., stride in Conv). Subsequent operators take the outputs of previous operators as their inputs, and the connection topology of these different operators constructs the model’s computational graph. The inputs and outputs of operators are four-dimensional tensors in [batch_num,channel,height,width][batch\_num,channel,height,width] format. During model inference, data is propagated through the computational graph until the final output is obtained. Quantization is a technique that reduces the precision of model weights. This process significantly decreases memory and computational requirements, enabling efficient inference on resource-limited edge devices, typically with minimal loss in accuracy.
DL compilers provide valuable solutions for deploying DNNs on edge devices. DL compilers perform a series of optimizations on these operators and ultimately compile them into executables. The operators are compiled into functions in executables, the inputs and ouputs of a operator correspond in sequence to the parameters of the operator function. The compiled executables can efficiently perform inference on resource-constrained edge devices. DL compilers also support quantized compiled models, which further reduces the resource consumption of model inference.
The general workflow of a DL compiler involves three steps: frontend processing, which converts general model representations, such as ONNX [33], into computational graphs supported by the compiler’s frontend; compilation optimization, applying various optimization techniques, including high-level optimizations like operator fusion and constant folding, and low-level optimizations such as layout rearrangement; code generation, which generates executables adapted for the target device’s hardware.
@@ II-C Research Gap
Works Optim Cross Arch Quant
@@
Libsteal cannot decompile standalone DNN executables and is unable to handle compilation optimizations. The accuracy of the models recovered by Libsteal is relatively low. The work by Shi et al. only supports x86 architecture and cannot recover models with high accuracy. DND and Neuroscope do not effectively handle compilation optimizations. DND relies on symbolic execution methods, which incur significant overhead and limit the size of supported models; Neuroscope only supports 12 DNN operators. Although BTD considers the impact of compilation optimizations, its operator type recognition method relies on prior knowledge like compiler versions, presenting challenges for extensibility, and it only supports x86 architecture. Moreover, all previous works overlook models compiled with quantization, which limits the applicability of these methods in practical scenarios.
@@
Previous DNN executables decompilers have their own limitations and existing decompilers struggle to simultaneously address compilation optimizations, support different architectures, and accommodate quantized compiled models. We summarize the challenges faced as follows. C1: A method for accurate operator type recognition needs to be designed. It should be able to cover a wide variety of operators and fused operators under compilation optimization. The method should not rely on prior knowledge such as compiler version or training data. C2: A decompiler needs to integrate the new method for operator type recognition, forming an end-to-end pipeline that recover DNN executables into high-level models. The decompiler should also have cross-architecture support capabilities. C3: Quantized compiled models exhibit new characteristics different from standard models, involving quantization scaling factors between integer and float domains. The decompiler needs to be compatible with these differences.
@@ III Approach
As shown in Figure 1, NeuroDeX is designed as an universal pipeline for different types of DNN executables, enabling end-to-end recovery of DNN executables into high-level models. Specifically, NeuroDeX first extracts operator function information, such as disassembled code, decompiled code and parameter dimensions from DNN executables. In this stage, NeuroDeX utilizes Ghidra [9], a general-purpose decompiler. Secondly, NeuroDeX completes operator type recognition and operator attribute recovery, where it leverages dynamic analysis and code semantic understanding from LLMs to support compatibility with various types of models. Dynamic analysis aims to monitor the runtime information of operator functions, and LLMs can understand the mathematical semantics of operators in decompiled code without relying on prior knowledge like compiler versions or training data. Finally, NeuroDeX performs model reconstruction by computational graph and model weights recovery, recovering high-level models. Next, we will sequentially introduce the components of NeuroDeX.
@@ III-A Operator Function Extraction
Inspired by previous works [59, 42], NeuroDeX can identify the dimensions of operator parameters from disassembled code in TVM compiler. NeuroDeX further expands on their methods, NeuroDeX also extracts the types of operator parameters and recover the optimized parameters’ dimensions. The parameters’ types can help determine the weights of quantized compiled models. The recovery of optimized parameters’ dimensions can help analyze optimized operators. Figure  2 is a specific example of Transform operator. The parameters’ dimensions and types can be extracted by scanning the disassembled code.
@@
The dimension of the first parameter is optimized into $[N,C,H,W,c]$ layout. NCHWc [3] is a commonly used inference optimization method, which adapts to hardware and accelerates inference. NeuroDeX can recover it into standard [N,Cc,H,W][N,C*c,H,W] format. Similarly, the dimension $O_{c},I_{c},K,K]$ of a Conv kernel parameter may be optimized into $O_{c}/A,I_{c}/B,K,K,B,A] [$3]. The optimized layout decides the storage order of model weights in memory. NeuroDeX extracts this optimized dimension layout and converts it to standard format. The parameters of the function correspond sequentially to the inputs and outputs of the operator. Each type operator has a fixed number of inputs and outputs, and NeuroDeX can accurately recover parameters’ dimensions. We validate this characteristic across both historical and recent TVM versions, ensuring the generality and robustness of our approach.
@@ III-B Operator Type Recognition
The purpose of operator type recognition is to determine the specific type of the operator function. Prior studies have limitations in operator type recognition. For instance, BTD relies on training on specific data of each compiler version, DND and Neuroscope perform bad in handling fused operators.
We manually analyze DL compiler’s support for DNN operators, and the operators can be divided into four types as shown in Table II. Type 1 is Layout Transformation. This type of operator adjusts the layout of input tensors and it can be further divided into inter-tensor types like Concat and intra-tensor types like Flatten. Inter-tensor types involve layout adjustments across multiple tensors, whereas intra-tensor refers to layout adjustments only to a single input tensor. Type 2 is Element-wise. This type of operator performs element-wise computation on the input tensor. It can be further divided into type 2.1 with two operands (one input and one output) like Softmax and type 2.2 with three operands (two inputs and one output) like Add. Type 3 is Reduction. This type of operator performs feature reduction on the input tensor, such as Maxpool. The output tensor is usually compressed in size compared to the input tensor. Type 4 is Complex. This type of operator performs complex and intensive computations, such as Conv. Due to the optimization of DL compilers, multiple operators may be fused into one operator to accelerate inference. TVM refers to enabling no optimization as “O0”, while enabling full optimizations as “O3”. GLOW applies full optimizations by default. Next, we will introduce the specific methods of operator type recognition towards TVM and GLOW.
@@
1. Layout Transformation 1.1 concat, split…
1.2 flatten, reshape, transpose…
@@
2. Element-wise 2.1 softmax, sqrt, clip, abs, relu…
2.2 add, sub, mul, div, power…
@@
3. Reduction 3.1 maxpool, avgpool, sum, max… Feature reduction and extraction
4. Complex 4.1 conv, dense… Complex and intensive computation
For TVM, as parameters’ dimensions of operator function can be collected during the operator function extraction phase, NeuroDeX uses them to enhance operator type recognition. Operators can undergo coarse-grained identification of their types. For Layout Transformation type and Complex type, specific operator types can be directly determined. For example, an operator with two parameters’ dimensions: $[1,1000,1,1]$ and $[1,1000]$ can be directly identified as Flatten. For Conv, input channel and output channel of kernel parameter correspond respectively to the channel of the first and last parameters. For Element-wise and Reduction type, operators can be determined candidate list among 2.1, 2.2, 3.1 in Table II according to the number and dimension of parameters. Further, leveraging the code semantic understanding capabilities of LLMs, NeuroDeX determines the specific operator types within the candidate list based on the mathematical features of the decompiled code. This approach overcomes the limitations of previous methods, which rely heavily on training data and struggle with the cross-version support for compilers. Here is the simplified prompt for specific operator type recognition:
@@
At the O3 optimization level, operators frequently undergo fusion. The most common operator fusion occurs Conv and Dense with the operators following them. NeuroDeX identifies their types through dynamic analysis. As shown in Algorithm 1, for Conv and Dense, during operator fusion, they are fused with several Multiply or Add operators, which are typically related to bias addition, batch normalization, etc. Specifically, NeuroDeX records the parameter addresses of each operator function through dynamic instrumentation. If the input is the output of previous operators, this indicates a skip connection (jumpadd). Otherwise, NeuroDeX starts recording the memory access during the operator function execution, identifying the instruction that initially accesses the parameter address. From this instruction, NeuroDeX performs taint analysis, tracking relevant registers until first encounter a multiply or add instruction. Activation functions like Relu and Clip are often attached to the tail of the fused operator with repeated patterns in decompiled code. NeuroDeX extracts the tail part of the decompiled code and uses LLM to verify if activation is accompanied by the fused operator. Similarly, for the fused operator consisting of Concat and activation, NeuroDeX employs the same processing method. NeuroDeX provides more comprehensive analysis and support for fused operators than previous works.
@@
For GLOW, the disassembled code does not contain dimension information, but the optimization strategy in GLOW is relatively simple. Operator fusion almost only occurs after Conv with activation. NeuroDeX directly performs operator type recognition by classifying the decompiled code with LLM, which considers all operators as the candidate list and determines the specific operator types within the candidate list based on the mathematical features of operators.
Currently, NeuroDeX supports the most common operators of ONNX, the complete list of operators can be found in Appendix IX-A. We compile and decompile validated computer vision models from the ONNX Model Zoo, and NeuroDeX can cover all the operators of them. The operator type recognition method of NeuroDeX has good scalability. For more operators, it only requires defining their types in Table II. The pipeline of operator type recognition in NeuroDeX is universal and scalable.
@@ III-C Operator Attribute Recovery
For Conv, NeuroDeX utilizes parameters’ dimensions of these operators to recover attributes. For TVM, dimensions have been collected from disassembled code introduced in operator function extraction. For GLOW, NeuroDeX employs dynamic analysis to collect parameters’ dimensions. The Conv in GLOW has four parameters: outout, inin, weightweight and biasbias. NeuroDeX records the memory read ranges starting from these four parameters. NeuroDeX first extracts kernel_size from decompiled code, the kernel_size is manifested as two continuous outer loop with identical values in decompiled code (An example can be found in Appendix IX-B). The output channel is equal to the number of weights read from the bias parameter. Next, the following relationships can be used:
@@
For Maxpool and Avgpool, NeuroDeX extracts channel value from decompiled code, which corresponds to the inner loops following two identical loops (An example can be found in Appendix IX-B). NeuroDeX records the ranges of input and output. Next, the following relationships can be used:
@@
Algorithm 2 explains how NeuroDeX accomplishes operator attribute recovery. For Conv or Pooling operators, the input_height, output_height and kernel_size are inherently contained within the parameters’ dimensions. The input_height $I_{h}$, padding $P$, kernel_size $K$, stride $S$, and output_height $O_{\text{h}}$ satisfy the following constraint:
@@
Local response normalization (lrn) has attributes: sizesize, $\beta$, $\alpha$, biasbias and Clip has attributes: minmin, maxmax. The attributes of lrn and clip generally have a large search space, making them unsuitable for simulation execution using dynamic analysis enumeration. However, these attributes are often obvious in the decompiled code, which can be extracted directly by LLM. Regarding Concat, the sequence of multiple inputs is an attribute that should be extracted. Moreover, at the O3 optimization level in TVM, Transpose may be fused after Concat. NeuroDeX instruments the executable to record input and output tensors of Concat. By enumerating the input order and the number of channel shuffle groups in Transpose, NeuroDeX simulates the forward of Concat until the computed tensor exactly matches the actual output tensor. Similarly, Transform may also involve Transpose fused afterward, NeuroDeX determines the attribute values same with Concat.
Here is the simplified prompt in operator attribute recovery by LLM:

### v2

@@ I Introduction
In recent years, deep learning (DL) has rapidly advanced in the real world. Deploying deep neural networks (DNNs) on edge devices can meet the real-time requirements of edge computing, enhance privacy protection and enable offline inference capabilities, making DNNs widely applicable in real-world scenarios. DL compilers, such as TVM [1] and GLOW [2], can compile high-level DNN models into executables for inference on edge devices. DNNs are composed of different neural network operators (e.g., Conv, Relu), and DL compilers compiles these operators into operator functions in executables. DL compilers optimize models during compilation to improve inference efficiency and reduce deployment environment dependencies, which provides a good solution for deploying models on edge devices [3, 4, 5].
In DNN executables, the operators and weights are compiled into incomprehensible machine code, thereby reducing the risk of model stealing attacks compared to white-box deployment. However, DNN executables may still pose security risks due to decompilation. DNN executables can potentially be decompiled to original high-level models. This undermines the intellectual property of the model owners, especially for models trained on private data. Based on the recovered high-level models, attackers can perform white-box adversarial attacks and backdoor attacks, threatening the secure use of DNN executables.
Previous works [6, 7, 8, 9, 10] have attempted to decompile DNN binaries, but the existing methods have certain limitations. They struggle to simultaneously address compilation optimizations, support different architectures, and accommodate quantized compiled model, all of which are essential for meeting the demands of modern DL applications. Previous decompilers also face issues such as insufficient accuracy and reliance on prior knowledge when recognizing operator types, and they also often come with significant analysis overhead.
@@ II-A Deep Learning Compiler
DNNs are computational models consisting of multiple layers. These layers are constructed from different types of neural network operators that perform various mathematical computations (e.g., Conv, Relu), and some operators have attributes (e.g., stride in Conv). Subsequent operators take the outputs of previous operators as their inputs, and the connection topology of these different operators constructs the model’s computational graph. The inputs and outputs of operators are four-dimensional tensors in [batch_num,channel,height,width][batch\_num,channel,height,width] format. During model inference, data is propagated through the computational graph until the final output is obtained. Quantization is a technique that reduces the precision of model weights. This process significantly decreases memory and computational requirements, enabling efficient inference on resource-limited edge devices, typically with minimal loss in accuracy.
DL compilers provide valuable solutions for deploying DNNs on edge devices. DL compilers perform a series of optimizations on these operators and ultimately compile them into executables. The operators are compiled into functions in executables, the inputs and ouputs of a operator correspond in sequence to the parameters of the operator function. The compiled executables can efficiently perform inference on resource-constrained edge devices. DL compilers also support quantized compiled models, which further reduces the resource consumption of model inference. The general workflow of a DL compiler involves three steps: frontend processing, which converts general model representations, such as ONNX [11], into computational graphs supported by the compiler’s frontend; compilation optimization, applying various optimization techniques, including high-level optimizations like operator fusion and constant folding, and low-level optimizations such as layout rearrangement; code generation, which generates executables adapted for the target device’s hardware.
@@ II-C Research Gap
Works Optimization Cross Arch Quantization
@@
Libsteal cannot decompile standalone DNN executables and is unable to handle compilation optimizations. The accuracy of the models recovered by Libsteal is relatively low. The work by Shi et al. only supports x86 architecture and cannot recover models with high accuracy. DND and Neuroscope do not effectively handle compilation optimizations. Although BTD considers the impact of compilation optimizations, it only supports x86 architecture. Moreover, all previous works overlook models compiled with quantization, which limits the applicability of these methods in practical scenarios. Previous DNN executables decompilers have their own limitations and existing decompilers struggle to simultaneously address compilation optimizations, support different architectures, and accommodate quantized compiled models.
Beyond the aforementioned discussion, we conduct a systematic analysis of operator type recognition, the critical step in decompilation pipeline. This analysis results in the following observations:
Observation1: Most previous works lack robust support for operator type recognition. Libsteal and Shi et al.’s work do not guarantee sufficient accuracy; DND relies on symbolic execution, which incurs significant overhead and limits the size of supported models; Neuroscope only supports 12 DNN operators. More importantly, these methods fail to provide adequate support for fused operators from compilation optimizations.
Observation2: BTD considers the impact of compilation optimizations and supports a wider range of operator types. However, BTD trains machine learning model for each compiler version to make predictions. This approach relies heavily on training data and treats the compiler version as prior knowledge, which limits its scalability in real world scenario. Moreover, we find that about 57.9% of the training data in TVM and 18.3% in GLOW appear in the test dataset, further undermining the effectiveness of the method. To avoid the influence of data leak, we randomly split all the data with an 8:2 ratio for train dataset and test dataset, and retrain the model for each compiler version, strictly following the default settings of BTD. As illustrated in Figure 1, we have analyzed BTD’s type recognition accuracy across different compiler versions. It is evident that BTD’s method exhibits poor cross-version support for compilers after avoiding data leak. Even within the same version, errors of operator type occur to an unacceptable degree, make it difficult to apply directly.
Figure 1: BTD Cross Version Type Recognition Accuracy
Observation3: To overcome the reliance on prior knowledge and improve recognition accuracy, we attempt to use a mathematical feature-based approach for operator type recognition. We specifically calculate the proportion of arithmetic operators (e.g., +, -, *, /) within the decompiled code of operator functions, and apply TSNE dimensionality reduction. The result is shown in Figure 2, the distributions of different types significantly overlap, indicating that classifying operators solely based on their mathematical characteristics is evidently challenging. Successfully performing operator type recognition requires a thorough capture of the semantic features of operator functions. The operator type recognition mechanism should have a good grasp of code semantic rather than strict mathematical feature matching.
Figure 2: TSNE Dimensionality Reduction of Different Operators
Reflecting on these observations, existing works struggle with operator type recognition. They exhibit limitations in the scope of covered operators and recognition accuracy.
@@
We summarize the challenges faced as follows. C1: A method for accurate operator type recognition needs to be designed. It should be able to cover a wide variety of operators and fused operators under compilation optimization. The method should not rely on prior knowledge such as compiler version or training data. C2: A decompiler needs to integrate the new method for operator type recognition, forming an end-to-end pipeline that recover DNN executables into high-level models. The decompiler should also have cross-architecture support capabilities. C3: Quantized compiled models exhibit new characteristics different from standard models, involving quantization scaling factors between integer and float domains. The decompiler needs to be compatible with these differences.
@@ III Approach
As shown in Figure 3, NeuroDeX is designed as an universal pipeline for different types of DNN executables, enabling end-to-end recovery of DNN executables into high-level models. Specifically, NeuroDeX first extracts operator function information, such as disassembled code, decompiled code and parameter dimensions from DNN executables. In this stage, NeuroDeX utilizes Ghidra [19], a general-purpose decompiler. Secondly, NeuroDeX completes operator type recognition and operator attribute recovery, where it leverages dynamic analysis and code semantic understanding from LLMs to support compatibility with various types of models. Dynamic analysis aims to monitor the runtime information of operator functions. Dynamic analysis in NeuroDeX requires only trivial input that satisfies the expected input format. This is due to the fact that any input can guarantee full coverage of the whole DNN model, and the mathematical dependencies of intermediate features are fixed. LLMs can understand the mathematical semantics of different operators in decompiled code without relying on prior knowledge like compiler versions or training data. Finally, NeuroDeX performs model reconstruction by computational graph and model weights recovery, recovering high-level models. Next, we will sequentially introduce the components of NeuroDeX.
@@ III-A Operator Function Extraction
Inspired by previous works [6, 9], NeuroDeX can identify the dimensions of operator parameters from disassembled code in TVM compiler. NeuroDeX further expands on their methods, NeuroDeX also extracts the types of operator parameters and recover the optimized parameters’ dimensions. The parameters’ types can help determine the weights of quantized compiled models. The recovery of optimized parameters’ dimensions can help analyze optimized operators. Figure 4 is a specific example of Transform operator. The parameters’ dimensions and types can be extracted by scanning the disassembled code.
@@
The dimension of the first parameter is optimized into $[N,C,H,W,c]$ layout. NCHWc [1] is a commonly used inference optimization method, which adapts to hardware and accelerates inference. NeuroDeX can recover it into standard [N,Cc,H,W][N,C*c,H,W] format. Similarly, the dimension $O_{c},I_{c},K,K]$ of a Conv kernel parameter may be optimized into $O_{c}/A,I_{c}/B,K,K,B,A] [$1]. The optimized layout decides the storage order of model weights in memory. NeuroDeX extracts this optimized dimension layout and converts it to standard format. This method is layout agnostic, it is compatible with other optimized layout, such as $[N,H,W,C]$. The parameters of the function correspond sequentially to the inputs and outputs of the operator. Each type operator has a fixed number of inputs and outputs, and NeuroDeX can accurately recover parameters’ dimensions. We validate this characteristic across both historical and recent TVM versions, ensuring the generality and robustness of our approach.
@@ III-B Operator Type Recognition
The purpose of operator type recognition is to determine the specific type of the operator function.
We manually analyze DL compiler’s support for DNN operators [1, 2] and align it with the practice of general DL frameworks like ONNX [11] to classify operators. The operators can be divided into four types as shown in Table II. Type 1 is Layout Transformation. This type of operator adjusts the layout of input tensors and it can be further divided into inter-tensor types like Concat and intra-tensor types like Flatten. Inter-tensor types involve layout adjustments across multiple tensors, whereas intra-tensor refers to layout adjustments only to a single input tensor. Type 2 is Element-wise. This type of operator performs element-wise computation on the input tensor. It can be further divided into type 2.1 with two operands (one input and one output) like Softmax and type 2.2 with three operands (two inputs and one output) like Add. Type 3 is Reduction. This type of operator performs feature reduction on the input tensor, such as Maxpool. The output tensor is usually compressed in size compared to the input tensor. Type 4 is Complex. This type of operator performs complex and intensive computations, such as Conv. Due to the optimization of DL compilers, multiple operators may be fused into one operator to accelerate inference. TVM refers to enabling no optimization as “O0”, while enabling full optimizations as “O3”. GLOW applies full optimizations by default. Next, we will introduce the specific methods of operator type recognition towards TVM and GLOW.
@@
1.Layout
Transformation
1.1 concat,split…
1.2 flatten,reshape,
transpose…
@@
2.Element-wise 2.1 softmax,sqrt,clip,
abs,relu…
2.2 add,sub,mul,div,
power…
@@
3.Reduction 3.1 maxpool,avgpool,
sum,max…
Feature reduction and extraction
4.Complex 4.1 conv,dense… Complex and intensive computation
For TVM, as parameters’ dimensions of operator function can be collected during the operator function extraction phase, NeuroDeX uses them to enhance operator type recognition. Operators can undergo coarse-grained identification of their types. For Layout Transformation type and Complex type, specific operator types can be directly determined. For example, an operator with two parameters’ dimensions: $[1,1000,1,1]$ and $[1,1000]$ can be directly identified as Flatten. For Conv, input channel and output channel of kernel parameter correspond respectively to the channel of the first and last parameters. For Element-wise and Reduction type, operators can be determined candidate list among 2.1, 2.2, 3.1 in Table II according to the number and dimension of parameters. Further, leveraging the code semantic understanding capabilities of LLMs, NeuroDeX determines the specific operator types within the candidate list based on the mathematical features of the decompiled code. Operator type recognition needs code semantic understanding rather than strict mathematical feature matching, LLMs are more suitable for it. This approach overcomes the limitations of previous methods, which rely heavily on training data and struggle with the cross-version support for compilers. NeuroDeX does not rely on complex prompt design, a simple and clear prompt can effectively accomplish the task, ensuring the stability of LLM participation in NeuroDeX. Here is the simplified prompt for specific operator type recognition:
@@
At the O3 optimization level, operators frequently undergo fusion. NeuroDeX identifies their types through dynamic analysis. The most common operator fusion occurs Conv and Dense with the operators following them. We use them as an example to illustrate how NeuroDeX handles fused operators. As shown in Algorithm 1, for Conv and Dense, during operator fusion, they are fused with several Multiply or Add operators, which are typically related to bias addition, batch normalization, etc. Specifically, NeuroDeX records the parameter addresses of each operator function through dynamic instrumentation. If the input is the output of previous operators, this indicates a skip connection (jumpadd). Otherwise, NeuroDeX starts recording the memory access during the operator function execution, identifying the instruction that initially accesses the parameter address. From this instruction, NeuroDeX performs taint analysis, tracking relevant registers until first encounter a multiply or add instruction. Activation functions like Relu and Clip are often attached to the tail of the fused operator with repeated patterns in decompiled code. NeuroDeX extracts the tail part of the decompiled code and uses LLM to verify if activation is accompanied by the fused operator. NeuroDeX employs the same processing method for other fused operators like the operator consisting of Concat and activation. NeuroDeX provides more comprehensive analysis and support for fused operators than previous works.
@@
For GLOW, the disassembled code does not contain dimension information, but the optimization strategy in GLOW is relatively simple. Operator fusion almost only occurs after Conv with activation. NeuroDeX directly performs operator type recognition by classifying the decompiled code with LLM, which considers all operators as the candidate list and determines the specific operator types within the candidate list based on the mathematical semantics of operators.
Currently, NeuroDeX supports the most common operators of ONNX (including but not limited to operators analyzed in previous works), the complete list of operators is shown below.
[…]
We compile and decompile validated computer vision models from the ONNX Model Zoo, and NeuroDeX can cover all the operators of them. The operator type recognition method of NeuroDeX has good scalability. For more operators, it only requires defining their types in Table II. The pipeline of operator type recognition in NeuroDeX is universal and scalable.
@@ III-C Operator Attribute Recovery
For Conv, NeuroDeX utilizes parameters’ dimensions of these operators to recover attributes. For TVM, dimensions have been collected from disassembled code introduced in operator function extraction. For GLOW, NeuroDeX employs dynamic analysis to collect parameters’ dimensions. The Conv in GLOW has four parameters: outout, inin, weightweight and biasbias. NeuroDeX records the memory read ranges starting from these four parameters. NeuroDeX first extracts kernel_size from decompiled code, the kernel_size is manifested as two continuous outer loop with identical values in decompiled code. The output channel is equal to the number of weights read from the bias parameter. Next, the following relationships can be used:
@@
For Maxpool and Avgpool, NeuroDeX extracts channel value from decompiled code, which corresponds to the inner loops following two identical loops. NeuroDeX records the ranges of input and output. Next, the following relationships can be used:
@@
The recovery of parameters’ dimensions in GLOW is also layout agnostic like TVM. NeuroDeX only needs the size of accessed memory region and does not concern the specific layout order. Algorithm 2 explains how NeuroDeX accomplishes operator attribute recovery. For Conv or Pooling operators, the input_height, output_height and kernel_size are inherently contained within the parameters’ dimensions. The input_height $I_{h}$, padding $P$, kernel_size $K$, stride $S$, and output_height $O_{\text{h}}$ satisfy the following constraint:
@@
Local response normalization (lrn) has attributes: sizesize, $\beta$, $\alpha$, biasbias and Clip has attributes: minmin, maxmax. The attributes of lrn and clip generally have a large search space, making them unsuitable for simulation execution using dynamic analysis enumeration. However, these attributes are often obvious in the decompiled code, which can be extracted directly by LLM. Regarding Concat, the sequence of multiple inputs is an attribute that should be extracted. Moreover, at the O3 optimization level in TVM, Transpose may be fused after Concat. NeuroDeX instruments the executable to record input and output tensors of Concat. By enumerating the input order and the number of channel shuffle groups in Transpose, NeuroDeX simulates the forward of Concat until the computed tensor exactly matches the actual output tensor. Similarly, Transform may also involve Transpose fused afterward, NeuroDeX determines the attribute values same with Concat. NeuroDeX only requires straightforward prompt for operator attribute recovery. Here is the simplified prompt in operator attribute recovery: